# Regularity of solutions to degenerate and singular free boundary problems with volume constraint

### v1

@@ 1. Introduction
$K_{0}=\big{\{}v\in H^{1}(B_{1},\omega):v\geq 0,\,v\big{|}_{\partial B_{1}}=g,\,\omega(\{v>0\}\cap B_{1})=m\big{\}}$
@@
Examples of $A_{2}$ weights for which (1.2) holds are $A_{1}$ weights, that is, weights $\omega$ for which there is a constant $C>0$ such that $M\omega(x)\leq C\omega(x)$ for a.e. $x\in\mathbb{R}^{n}$, where $M$ is the classical Hardy–Littlewood maximal function (note that $A_{1}\subset A_{2}$). Indeed, this is a consequence of the results in [Turesson00, Section 2.4.2]. In fact, (1.2) is stated as [Turesson00, Theorem 2.4.8] (this result is valid for our case because Theorem 2.4.7–the co-area formula, and Lemma 2.4.1–the Poincaré inequality, in [Turesson00] are true under our hypotheses as well). Furthermore, we do not require the $A_{1}$ weight to be continuous. Canonical examples of $A_{2}$ weights satisfying (1.3) are $\omega(x)=|x|^{-\alpha}$ for $0<\alpha<n$, or, more generally, $\omega(x)=\operatorname{dist}(x,Z)^{\alpha}$, where $Z$ is a “thin” set of zero Lebesgue measure, and anisotropic variations of power weights (see [LST] where the lower bound in $\omega$ is assumed to study a weighted one phase problem generalizing [ALtCaff81]). In particular, for the case of the extension problem for the fractional Laplacian $(-\Delta)^{s}$, where $\omega(x,y)=|y|^{1-2s}$ for $(x,y)\in\mathbb{R}^{n+1}$, we have the restriction $1/2\leq s<1$.
@@
for some parameter $\varepsilon>0$. Assuming only the $A_{2}$ condition on the weight, we show that minimizers of the penalized problem exist for all $\varepsilon>0$ and are Hölder continuous. The Hölder continuity is proved using a blow-up approximation by constants in the spirit of Caffarelli’s regularity technique. Notice that this technique was also implemented in [LST] for the one-phase problem. However, unlike [LST], we are able to establish Hölder continuity of solutions to our problem (1.1) without imposing any further restrictions on the weight $\omega$ (other than being in the $A_{2}$ class). Then we show that there exists $\varepsilon_{\ast}>0$ small enough such that $u_{\ast}$ is a minimizer for the penalized problem if and only if $u_{\ast}$ is a minimizer for (1.1). Here we argue by contradiction, finding an appropriate competitor function $u_{t}$ for which $\{u_{t}>0\}$ approximates in measure $\{u>0\}$, and we apply the hypothesis on $\omega$ stated in Theorem 1.1.
@@ 2. Preliminaries
By definition of $A_{2}$ weights, we may assume that $0<\omega<\infty$ almost everywhere. Thus, $\omega$ defines a Radon measure given by $\omega(E):=\int_{E}\omega(x)\,dx$. Moreover, the weight $\tilde{\omega}(x):=\omega(x_{0}+rx)$, for fixed $x_{0}\in\mathbb{R}^{n}$ and $r>0$, is also an $A_{2}$ weight with $\tilde{\omega}]_{A_{2}}=[\omega]_{A_{2}}$.
@@ 3. The penalized problem with parameter $\varepsilon>0$
where $C>0$ is independent of $j$. Thus, there exists $u\in H^{1}(B_{1},\omega)$ such that, up to a subsequence, $v_{j}\to u$ weakly in $H^{1}(B_{1},\omega)$. By the compact embedding $H^{1}(B_{1},\omega)\subset\subset L^{2}(B_{1},\omega)$ and (2.2), up to a further subsequence, $v_{j}\to u$ strongly in $L^{2}(B_{1},\omega)$ and a.e. in $B_{1}$. In particular, $u\geq 0$ a.e.. Moreover, since $v_{j}-g$ weakly converges in $H^{1}_{0}(B_{1},\omega)$, by Mazur’s lemma, it follows that $u-g\in H^{1}_{0}(B_{1},\omega)$, so that $u\in H^{1}_{g}(B_{1},\omega)$.
@@
Note that the term $\omega(\{u>0\}\cap(B\setminus E))$ is a constant in the minimization in Definition 3.3 and that a minimizer $u$ of $J_{\varepsilon}(u,\omega,B_{1})$ is a local minimizer in $B_{1}$. For $x_{0}\in B_{1/2}$, $0<r<1/2$ and $\delta>0$, we define the rescaled function
@@ 4. Behavior for small $\varepsilon$
To estimate this term, recall that $u\in W^{1,1}(B_{1})$ is locally Hölder continuous by Theorem 3.5. Then we can apply the co-area formula for Sobolev functions from [maly-swanson-ziemer] to get

### v2

@@ 1. Introduction
$K_{0}=\big\{v\in H^{1}(B_{1},\omega):\,v\big|_{\partial B_{1}}=g,\,\omega(\{v>0\}\cap B_{1})=m\big\}$
@@
Examples of $A_{2}$ weights for which (1.2) holds are $A_{1}$ weights, that is, weights $\omega$ for which there is a constant $C>0$ such that $M\omega(x)\leq C\omega(x)$ for a.e. $x\in\mathbb{R}^{n}$, where $M$ is the classical Hardy–Littlewood maximal function (note that $A_{1}\subset A_{2}$). Indeed, this is a consequence of the results in [Turesson00, Section 2.4.2]. In fact, (1.2) is stated as [Turesson00, Theorem 2.4.8] (this result is valid for our case because Theorem 2.4.7–the co-area formula, and Lemma 2.4.1–the Poincaré inequality, in [Turesson00] are true under our hypotheses as well). Furthermore, we do not require the $A_{1}$ weight to be continuous. Canonical examples of $A_{2}$ weights satisfying (1.3) are $\omega(x)=|x|^{-\alpha}$ for $0<\alpha<n$, or, more generally, $\omega(x)=\operatorname{dist}(x,Z)^{-\alpha}$, where $Z$ is a “thin” set of zero Lebesgue measure, and anisotropic variations of power weights (see [LST] where the lower bound in $\omega$ is assumed to study a weighted one phase problem generalizing [ALtCaff81]). In particular, for the case of the extension problem for the fractional Laplacian $(-\Delta)^{s}$, where $\omega(x,y)=|y|^{1-2s}$ for $(x,y)\in\mathbb{R}^{n+1}$, we have the restriction $1/2\leq s<1$.
@@
for some parameter $\varepsilon>0$. Assuming only the $A_{2}$ condition on the weight, we show that minimizers of the penalized problem exist for all $\varepsilon>0$ and are Hölder continuous. The Hölder continuity is proved using a blow-up approximation by constants in the spirit of Caffarelli’s regularity technique. Notice that this technique was also implemented in [LST] for the one-phase problem. However, unlike [LST], we are able to establish Hölder continuity of solutions to our problem (1.1) without imposing any further restrictions on the weight $\omega$ (other than being in the $A_{2}$ class). Then we show that there exists $\varepsilon_{\ast}>0$ small enough such that $u_{\ast}$ is a minimizer for the penalized problem if and only if $u_{\ast}$ is a minimizer for (1.1). Here we argue by contradiction, finding an appropriate competitor function $u_{t}$ for which $\{u_{t}>0\}$ approximates in measure $\{u>0\}$, and we apply the hypothesis on $\omega$ stated in Theorem 1.1. This argument for proving existence of minimizers was implemented in [BHP05] in the unweighted case, where Lipschitz continuity of minimizers was also proved.
@@
An interesting question is whether existence of minimizers fails if one removes either assumption (1.2) or (1.3). We believe that one should be able to remove them, but the technique will have to be different.
@@ 2. Preliminaries
By definition of $A_{2}$ weights, we may assume that $0<\omega<\infty$ almost everywhere. Moreover, as $\omega\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{n})$, the induced measure defined by
$\omega(E):=\int_{E}\omega(x)\,dx$
is a positive Radon measure. Moreover, the weight $\tilde{\omega}(x):=\omega(x_{0}+rx)$, for fixed $x_{0}\in\mathbb{R}^{n}$ and $r>0$, is also an $A_{2}$ weight with $\tilde{\omega}]_{A_{2}}=[\omega]_{A_{2}}$.
@@ 3. The penalized problem with parameter $\varepsilon>0$
where $C>0$ is independent of $j$. Thus, there exists $u\in H^{1}(B_{1},\omega)$ such that, up to a subsequence, $v_{j}\to u$ weakly in $H^{1}(B_{1},\omega)$. By the compact embedding $H^{1}(B_{1},\omega)\subset\subset L^{2}(B_{1},\omega)$ and (2.2), up to a further subsequence, $v_{j}\to u$ strongly in $L^{2}(B_{1},\omega)$ and a.e. in $B_{1}$. In particular, $u\geq 0$ a.e.. Moreover, since $v_{j}-g$ weakly converges in $H^{1}_{0}(B_{1},\omega)$, and $H^{1}_{0}(B_{1},\omega)$ is convex and strongly closed, it follows that $u-g\in H^{1}_{0}(B_{1},\omega)$, so that $u\in H^{1}_{g}(B_{1},\omega)$.
@@
Note that the term $\omega(\{u>0\}\cap(B\setminus E))$ is a constant in the minimization in Definition 3.3 and that a minimizer $u$ of $J_{\varepsilon}(u,\omega,B_{1})$ is a local minimizer in $B_{1}$. Indeed, if there were a set $E$ and a function $v$ on $E$ for which the above functional is smaller than for $u$, then one could take $w=u$ in $B\setminus E$ and $w=v$ in $E$ to get $J_{\varepsilon}(w,\omega,B_{1})<J_{\varepsilon}(u,\omega,B_{1})$. By taking $E=B_{1}$, any local minimizer is a (global) minimizer, however it may not satisfy our imposed boundary condition so we do not use this latter property.
For $x_{0}\in B_{1/2}$, $0<r<1/2$ and $\delta>0$, we define the rescaled function
@@ 4. Behavior for small $\varepsilon$
To estimate this term, recall that $u\in W^{1,1}(B_{1})$ is locally Hölder continuous by Theorem 3.5. Then we can apply the co-area formula for Sobolev functions from [maly-swanson-ziemer, Theorem 1.4] to get